OpenStudy (anonymous):

Which of the following is the slope between the points (3, -2) and (8, 4)?

OpenStudy (anonymous):

it's 6/5

OpenStudy (lgbasallote):

\[\frac{y_2 - y_1}{x_2 - x_1}\] \[\implies \frac{4-(-2)}{8-3}\] \[\implies \frac{4+2}{8-3}\] \[\implies \frac 65\] you're right